OpenStudy (anonymous):

solve: y divided by -6+5=9

ganeshie8 (ganeshie8):

\(\large \frac{y}{-6} + 5 = 9\)

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

or is it y/(6+5)=9

OpenStudy (anonymous):

i mean -6 sorry

OpenStudy (anonymous):

the first one at the top

ganeshie8 (ganeshie8):

ok, we subtract 5

ganeshie8 (ganeshie8):

\(\large \frac{y}{-6} + 5 = 9 \) -5 -5

ganeshie8 (ganeshie8):

\(\large \frac{y}{-6} = 4\)

ganeshie8 (ganeshie8):

multiply -6

ganeshie8 (ganeshie8):

\(\large \frac{y}{-6} = 4\) x(-6) x(-6)

ganeshie8 (ganeshie8):

\(y = -24\)

OpenStudy (anonymous):

thank youu

ganeshie8 (ganeshie8):

you're wlcme :)